# A pipeline for enabling path-specific causal fairness in observational health data

### v1

@@ Abstract.
When training machine learning (ML) models for potential deployment in a healthcare setting, it is essential to ensure that they do not replicate or exacerbate existing healthcare biases. Although many definitions of fairness exist, we focus on path-specific causal fairness, which allows us to better consider the social and medical contexts in which biases occur (e.g., direct discrimination by a clinician or model versus bias due to differential access to the healthcare system) and to characterize how these biases may appear in learned models. In this work, we map the structural fairness model to the observational healthcare setting and create a generalizable pipeline for training causally fair models [^1] Code available at https://github.com/reAIM-Lab/causal_fairness_health_pipeline. The pipeline explicitly considers specific healthcare context and disparities to define a target ”fair” model. Our work fills two major gaps: first, we expand on characterizations of the ”fairness-accuracy” tradeoff by detangling direct and indirect sources of bias and jointly presenting these fairness considerations alongside considerations of accuracy in the context of broadly known biases. Second, we demonstrate how a foundation model trained without fairness constraints on observational health data can be leveraged to generate causally fair downstream predictions in tasks with known social and medical disparities. This work presents a model-agnostic pipeline for training causally fair machine learning models that address both direct and indirect forms of healthcare bias.
@@
Causal fairness, healthcare, foundation models, causal inference, observational health data, fair machine learning
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 27–30, 2026; Montreal, CAisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Health informatics
1. Introduction
@@
The need for fairness in machine learning (ML) is well-documented, particularly in the context of healthcare, where there are known systemtic disparities that can lead to bias in data collection and in downstream ML models (Chen et al., 2021). Bias in the healthcare system exists through several avenues: implicit biases held by clinicians can directly the lower quality of patient care (FitzGerald and Hurst, 2017), and inequitable exposures to social and environmental factors (e.g., environmental toxicants) can indirectly cause disparities in health outcomes (Bellavia et al., 2018). Additionally, disparities in access and trust in the healthcare system can lead to disparities in healthcare utilization (Manuel, 2018). Differences in healthcare utilization can impact the completeness of observational health data in particular, as these data are only collected when individuals access healthcare services (Sullivan et al., 2019). We argue that, when training and evaluating data-driven models for potential use in healthcare settings, these different ”avenues” of bias must be characterized in a case-specific manner that accounts for known healthcare disparities in the specific disease domain.
@@
Many notions of fairness exist in machine learning (Gao et al., 2025). Group fairness is defined by splitting a population into groups based on some sensitive attribute(s) and enforcing equal model performance across subgroups according to some metric(s). However, notions of group fairness are limited by assumptions about the underlying data (e.g., demographic parity assumes equal rates of positive outcomes across all groups) and by the conflicting nature of group fairness metrics that can render simultaneous enforcement impossible (Chouldechova, 2017). Individual fairness defines a similarity metric and seeks to ensure that individuals defined as ”similar” by this metric have similar outcomes under the model (Dwork et al., 2012). However, individual fairness is susceptible to biased data and can overlook difficult-to-capture socioeconomic privileges and disparities in resource allocation (Braveman, 2014). Causal fairness looks to identify and mitigate ”unfair” causal pathways between some characteristic(s) and the outcome. Causal fairness approaches allow users to compare bias in the data to bias in the model, which is not possible when evaluating disparities based on subgroup comparison of model performance metrics. However, causal fairness can be limited by the need for domain knowledge to identify the underlying causal structure of the data and the need for statistical identifiability to estimate causal effects from observational data (Pearl, 2000; Makhlouf et al., 2024). Additionally, calculating counterfactual quantities in high dimensions can be complex and unstable  (Mitra et al., 2022).
@@
Figure 1. Standard Fairness Model: (a) Overview of SFM as presented in Plečko and Bareinboim (2024) (b) Clinically motivated mapping of SFM for observational health data
@@
Causal directed acyclic graphs illustrating the standard fairness model and its mapping to observational health data, showing relationships between sensitive attributes, clinical features, healthcare utilization, and outcome
@@
We argue that evaluation of fairness in healthcare should rely on path-specific causal fairness, which defines biased causal paths in the data-generating mechanism and seeks to ensure that the model does not replicate or introduce bias through these ”unfair” causal pathways. We focus on path-specific causal fairness over other methods, such as counterfactual fairness (Kusner et al., 2017) because path-specific causal fairness allows us to decompose the biases in the healthcare system and form hypotheses about specific social and medical disparities that a model may be replicating. Our work builds on the structural fairness model (SFM) defined in Plečko and Bareinboim (2024), which allows us to cluster variables into four groups (Figure 1a): sensitive attribute ($X$), outcome ($Y$), confounding variables ($Z$), and mediators ($W$). Through this clustering, we can ignore the relationships between variables within each cluster while maintaining identifiability of the relevant causal effects (Anand et al., 2023). We offer a generalizable mapping of high-dimensional observational health data onto the SFM (Section 3.2, Figure 1b), eliminating the need for domain knowledge of the case-specific specific causal system.
@@
We establish a pipeline for enforcing and evaluating path-specific causal fairness in clinical risk prediction models. Users can select target effects for fairness interventions and explore a more multidimensional fairness-accuracy tradeoff that decomposes ”bias” into specific paths that echo known health disparities. We demonstrate the pipeline’s generalizability by applying it to four case studies. This generalizability is important given the rise of structured electronic health record (EHR) foundation models for healthcare (Pang et al., 2025), which require adaptable methods for models that may not be trained to explicitly address fairness. Three of our cases (acute myocardial infarction, systemic lupus erythematosus, type 2 diabetes mellitus) leverage embeddings from a structured EHR foundation model (Wornow et al., 2025) to train the baseline model on which fairness interventions are applied. Our fourth case (schizophrenia risk prediction) is built on a task-specific model to demonstrate the utility of our pipeline on a more ”conventional” ML model.
@@
  3. (3)
    We establish a pipeline for ”diagnosing”, enforcing, and evaluating path-specific causal fairness by contextualizing data-driven insights with known. healthcare disparities
  4. (4)
@@
2. Related Work
2.1. Enforcing causal fairness in machine learning
@@
Training causally fair models is an open challenge in machine learning. To train counterfactually fair models, Kusner et al. (2017) propose limiting a model to learn only from features that are non-descendants of the sensitive attribute; however, this technique is not feasible in observational health data, where most clinical variables could be causal descendants of the sensitive attribute  (Gopal et al., 2021; Meidert et al., 2023; Vohra-Gupta et al., 2023). Counterfactual fairness is also enforced using learned representations that are invariant or orthogonal to the sensitive attribute (Quinzan et al., 2024; Chen and Zhu, 2025) or through counterfactual data augmentation (Ma et al., 2023; Zhou et al., 2025; Robertson et al., 2025).
@@
To enforce causal fairness with respect to path-specific measures, Plečko and Bareinboim (2024) present an in-processing regularization to penalize causal effects along specific paths in the structural fairness model (SFM) and enforce fair decision-making  (Plecko and Bareinboim, 2023). Other work develops interventions to reduce only the direct effect (Stefano et al., 2020) or the effect of spurious features (Wang et al., 2021b). Optimal transport has also been used to learn causally fair models (De Lara et al., 2024; Bayer et al., 2023; Nabi and Shpitser, 2018). In particular, Nabi et al. (2019) propose to learn fair policies by sampling data from a ‘debiased’ data-generating process. A full review of causality-based methods for fair ML can be found in Su et al. (2022). We compare path-specific in-processing and causally fair resampling to naive feature selection methods for their potential to reduce path-specific effects and find that methods specifically designed to address causal fairness do not universally outperform naive methods, even when fairness is evaluated using path-specific causal effects.
2.2. Navigating the ”fairness-accuracy tradeoff”
@@
Fairness and accuracy are important but potentially conflicting desiderata within ML (Gittens et al., 2022), with many works conceptualizing this tradeoff as a Pareto frontier (Valdivia et al., 2021; Wang et al., 2021a; Xu and Strohmer, 2023). Prior work in causality has aimed to measure this tradeoff, including by quantifying the extent to which model loss increases when trained under causal fairness constraints (Plečko and Bareinboim, 2025) and by using the ”average treatment effect” to assess the impact of fairness interventions (Ji et al., 2024). However, the assumption that fairness and accuracy must exist as a ”tradeoff” may be due to a lack of consideration of social and historical context (Wang et al., 2021a), limitations of the mathematical notations used to model fairness, and measurement error in biased data (Dutta et al., 2020). Our work explicitly considers social context and baseline bias in the observational data by accounting for non-direct causal effects, as observational health data is known to reflect the biases of our medical system. For a more in-depth treatment of the necessity of the fairness-accuracy tradeoff, we refer readers to (Li and Li, 2025). We emphasize the importance of using path-specifc causal fairness to assess ML models in the healthcare setting, breaking down bias into direct, indirect, and spurious causal effects. This allows us to consider the model in the context of braoder health disparities, quantify the extent to which a learned model replicates and/or exacerbates unfair causal paths in the data, and measure the impact of fairness interventions on model performance.
2.3. Fairness in foundation models
@@
Most work around fairness in foundation models focuses on language models (Wang et al., 2025; Gallegos et al., 2024) and vision models (Ali et al., 2023b). In medical imaging specifically, prior work found significant subgroup performance differences in segmentation foundation models (Li et al., 2024; Jin et al., 2024). Other work echoes these findings and addresses these differences using balanced fine-tuning datasets (Khan et al., 2023) or ”fair PCA post-processing” (Ali et al., 2023a). Consistent with broader ethical ML literature (Chen et al., 2021), Queiroz et al. (2025) discusses the ways that bias mitigation for medical vision foundation models must be an integrated effort involving investment at all steps in the AI pipeline, from data collection to training methodologies to deployment and regulation. Relatively little work focuses on fairness in tabular foundation models. Robertson et al. (2025) demonstrates how training on causally fair synthetic data can improve the fairness of tabular foundation models. In the case of in-context learning, Kenfack et al. (2025) illustrates the promise of pre-processing fairness methods. Our work examines the path-specific causal fairness of structured (tabular) EHR foundation models, both at the pretraining step and after linear probing, which has received little attention in prior work.
3. Preliminaries
3.1. Standard Fairness Model
@@
The standard fairness model (SFM) (Plečko and Bareinboim, 2024) is a structural causal model (SCM) to assess causal fairness. An SCM (Pearl, 2000) is a formalized model of a data generating process that defines the causal relationship between each of the variables in the system. The SFM defines the sensitive attribute ($X$) that can directly impact the outcome ($Y$). Additionally, $X$ affects $Y$ through a set of mediator variables ($W$), confounded by $Z$ (Figure 1a). One key aspect of the SFM is the identifiability of the natural direct effect (NDE), natural indirect effect (NIE), and spurious effect (SE) from observational data. That is, under the assumption that the data-generating mechanism follows the SFM, all three effects can be estimated as nested counterfactuals from observational data (Appendix A).
3.2. Clinical mapping of Standard Fairness Model
@@
We leverage the SFM to define a generalizable procedure for mapping high-dimensional observational health data (e.g., EHRs, administrative claims) onto the SFM, thus providing a principled template for evaluating fairness in observational health settings. Such a generalizable mapping is important because relying on clinical domain knowledge to define structural assumptions is infeasible in high dimensions. Our mapping (Figure 1b) is as follows: as with the SFM, $X$ remains the sensitive attribute and $Y$ is the clinical outcome of interest. $W$ represents a high-dimensional representation of a person’s clinical record that can include information derived from diagnoses, medications, procedures, devices, laboratory tests, etc. These features can be derived directly from a person’s clinical record or can be a learned representation of the clinical data presented as a set of embedding vectors. Finally, $Z$ represents other demographic variables, baseline comorbidities known before a visit, if available, and healthcare utilization. Healthcare utilization (HCU) refers to the frequency, types, and patterns of healthcare services that an individual uses. HCU can serve as a proxy for several unobserved variables such as social determinants of health (typically constituting latent common causes between $X$ and $Z$), including access to and trust in the healthcare system. In our work, we operationalize healthcare utilization as the frequency of visits. The NDE represents the ”direct” effect of the sensitive attribute on the outcome; the NIE represents the effect of the sensitive attribute on the outcome as mediated by the rest of a person’s observational health record ($W$). The SE represents the baseline change in the outcome distribution as a result of the two interventions $x=0$ and $x=1$ compared to their observational counterparts.
4. Methods
@@
Figure 2. Overview of the proposed workflow. Starting with task selection, we outline the steps required to train and evaluate a causally fair model. In particular, we find that data-driven decision making must be combined with known historical context about healthcare disparities specific to the clinical task at hand.
@@
Flowchart presenting the workflow outlined in this paper. The steps are: 1) define the task and key variables (in particular, select the sensitive attribute and calculate healthcare utilization); 2) robustly estimate the path-specific causal effects; 3) train a baseline model; 4) select target effects (e.g., NDE, NIE, SE, some combination) for which to reduce disparities; 5) perform and evaluate causal fairness enhancing interventions.
@@
4.1. Define the task and key variables
Once a clinical prediction task is selected, we use this to set our outcome ($Y$) in the SFM. The sensitive attribute ($X$), however, still requires definition. $X$ is typically a discrete variable, often binary, for which we want to measure and enforce path-specific fairness. While purely data-driven analysis based on all known sensitive attributes is important, it is often more crucial and feasible to assess how broadly known disparities reflect in the data and downstream learned models. We advocate for selecting a sensitive attribute that is associated with known (or hypothesized) disparities in broader health literature. We use race or gender as our main sensitive attributes for each case; however, selecting subgroups for $X$ using an intersection of identities is possible under statistical feasibility (sufficient sample sizes to robustly evaluate path-specific effects). The next key variable(s), often ignored, are related to healthcare utilization. Healthcare utilization patterns drive predictive performance of learned models  (Zink et al., 2024; Pang et al., 2025) and serve as noisy proxies for unobserved social drivers of health. We define healthcare utilization as the number of visits per year. If there is consistent information about the visit setting (e.g., outpatient, inpatient, pharmacy), we create separate healthcare utilization features for each visit setting. These healthcare utilization features, in combination with any ”non-sensitive” demographic features, comprise the confounder set of the SFM ($Z$). $Z$ must also include baseline health status, such as known comorbidities known at the beginning of the visit or stay, if data samples represent visits or stays. The clinical features ($W$) include all features used by the model that are not encapsulated by $X$, $Z$, and $Y$.
4.2. Robustly estimate path-specific causal effects

### v2

@@ Abstract
When training machine learning (ML) models for potential deployment in a healthcare setting, it is essential to ensure that they do not replicate or exacerbate existing healthcare biases. Although many definitions of fairness exist, we focus on path-specific causal fairness, which allows us to better consider the social and medical contexts in which biases occur (e.g., direct discrimination by a clinician or model versus bias due to differential access to the healthcare system) and to characterize how these biases may appear in learned models. In this work, we map the structural fairness model to the observational healthcare setting and create a generalizable pipeline for training causally fair models [^1] Code available at https://github.com/reAIM-Lab/causal_fairness_health_pipeline. The pipeline explicitly considers specific healthcare context and disparities to define a target "fair" model. Our work fills two major gaps: first, we expand on characterizations of the "fairness-accuracy" tradeoff by detangling direct and indirect sources of bias and jointly presenting these fairness considerations alongside considerations of accuracy in the context of broadly known biases. Second, we demonstrate how a foundation model trained without fairness constraints on observational health data can be leveraged to generate causally fair downstream predictions in tasks with known social and medical disparities. This work presents a model-agnostic pipeline for training causally fair machine learning models that address both direct and indirect forms of healthcare bias.
@@
Keywords Causal fairness $\cdot$ foundation models $\cdot$ causal inference $\cdot$ observational health data $\cdot$ fair machine learning
1 Introduction
@@
The need for fairness in machine learning (ML) is well-documented, particularly in the context of healthcare, where there are known systemic disparities that can lead to bias in data collection and in downstream ML models (Chen et al., 2021). Bias in the healthcare system exists through several avenues: implicit biases held by clinicians can directly the lower quality of patient care (FitzGerald and Hurst, 2017), and inequitable exposures to social and environmental factors (e.g., environmental toxicants) can indirectly cause disparities in health outcomes (Bellavia et al., 2018). Additionally, disparities in access and trust in the healthcare system can lead to disparities in healthcare utilization (Manuel, 2018). Differences in healthcare utilization can impact the completeness of observational health data in particular, as these data are only collected when individuals access healthcare services (Sullivan et al., 2019). We argue that, when training and evaluating data-driven models for potential use in healthcare settings, these different "avenues" of bias must be characterized in a case-specific manner that accounts for known healthcare disparities in the specific disease domain.
@@
Many notions of fairness exist in machine learning (Gao et al., 2025). Group fairness is defined by splitting a population into groups based on some sensitive attribute(s) and enforcing equal model performance across subgroups according to some metric(s). However, notions of group fairness are limited by assumptions about the underlying data (e.g., demographic parity assumes equal rates of positive outcomes across all groups) and by the conflicting nature of group fairness metrics that can render simultaneous enforcement impossible (Chouldechova, 2017). Individual fairness defines a similarity metric and seeks to ensure that individuals defined as "similar" by this metric have similar outcomes under the model Dwork et al. (2012). However, individual fairness is susceptible to biased data and can overlook difficult-to-capture socioeconomic privileges and disparities in resource allocation (Braveman, 2014). Causal fairness looks to identify and mitigate "unfair" causal pathways between some characteristic(s) and the outcome. Causal fairness approaches allow users to compare bias in the data to bias in the model, which is not possible when evaluating disparities based on subgroup comparison of model performance metrics. However, causal fairness can be limited by the need for domain knowledge to identify the underlying causal structure of the data and the need for statistical identifiability to estimate causal effects from observational data (Pearl, 2000; Makhlouf et al., 2024). Additionally, calculating counterfactual quantities in high dimensions can be complex and unstable  (Mitra et al., 2022).
@@
Figure 1: Standard Fairness Model: (a) Overview of SFM as presented in Plečko and Bareinboim (2024) (b) Clinically motivated mapping of SFM for observational health data
@@
Figure 2: Causal directed acyclic graphs illustrating the standard fairness model and its mapping to observational health data, showing relationships between sensitive attributes, clinical features, healthcare utilization, and outcome
@@
We argue that evaluation of fairness in healthcare should rely on path-specific causal fairness, which defines biased causal paths in the data-generating mechanism and seeks to ensure that the model does not replicate or introduce bias through these "unfair" causal pathways. We focus on path-specific causal fairness over other methods, such as counterfactual fairness Kusner et al. (2017) because path-specific causal fairness allows us to decompose the biases in the healthcare system and form hypotheses about specific social and medical disparities that a model may be replicating. Our work builds on the structural fairness model (SFM) defined in Plečko and Bareinboim (2024), which allows us to cluster variables into four groups (Figure 2a): sensitive attribute ($X$), outcome ($Y$), confounding variables ($Z$), and mediators ($W$). Through this clustering, we can ignore the relationships between variables within each cluster while maintaining identifiability of the relevant causal effects (Anand et al., 2023). We offer a generalizable mapping of high-dimensional observational health data onto the SFM (Section 3.2, Figure 2b), eliminating the need for domain knowledge of the case-specific causal system.
@@
We establish a pipeline for enforcing and evaluating path-specific causal fairness in clinical risk prediction models. Users can select target effects for fairness interventions and explore a more multidimensional fairness-accuracy tradeoff that decomposes "bias" into specific paths that echo known health disparities. We demonstrate the pipeline’s generalizability by applying it to four case studies. This generalizability is important given the rise of structured electronic health record (EHR) foundation models for healthcare Pang et al. (2025), which require adaptable methods for models that may not be trained to explicitly address fairness. Three of our cases (acute myocardial infarction, systemic lupus erythematosus, type 2 diabetes mellitus) leverage embeddings from a structured EHR foundation model Wornow et al. (2025) to train the baseline model on which fairness interventions are applied. Our fourth case (schizophrenia risk prediction) is built on a task-specific model to demonstrate the utility of our pipeline on a more "conventional" ML model.
@@
  3. 3.
    We establish a pipeline for "diagnosing", enforcing, and evaluating path-specific causal fairness by contextualizing data-driven insights with known. healthcare disparities
  4. 4.
@@
2 Related Work
2.1 Enforcing causal fairness in machine learning
@@
Training causally fair models is an open challenge in machine learning. To train counterfactually fair models, Kusner et al. (2017) propose limiting a model to learn only from features that are non-descendants of the sensitive attribute; however, this technique is not feasible in observational health data, where most clinical variables could be causal descendants of the sensitive attribute  (Gopal et al., 2021; Meidert et al., 2023; Vohra-Gupta et al., 2023). Counterfactual fairness is also enforced using learned representations that are invariant or orthogonal to the sensitive attribute Quinzan et al. (2024); Chen and Zhu (2025) or through counterfactual data augmentation (Ma et al., 2023; Zhou et al., 2025; Robertson et al., 2025).
@@
To enforce causal fairness with respect to path-specific measures, Plečko and Bareinboim (2024) present an in-processing regularization to penalize causal effects along specific paths in the structural fairness model (SFM) and enforce fair decision-making  (Plecko and Bareinboim, 2023). Other work develops interventions to reduce only the direct effect Stefano et al. (2020) or the effect of spurious features (Wang et al., 2021b). Optimal transport has also been used to learn causally fair models (De Lara et al., 2024; Bayer et al., 2023; Nabi and Shpitser, 2018). In particular, Nabi et al. (2019) propose to learn fair policies by sampling data from a ‘debiased’ data-generating process. A full review of causality-based methods for fair ML can be found in Su et al. (2022). We compare path-specific in-processing and causally fair resampling to naive feature selection methods for their potential to reduce path-specific effects and find that methods specifically designed to address causal fairness do not universally outperform naive methods, even when fairness is evaluated using path-specific causal effects.
2.2 Navigating the "fairness-accuracy tradeoff"
@@
Fairness and accuracy are important but potentially conflicting desiderata within ML Gittens et al. (2022), with many works conceptualizing this tradeoff as a Pareto frontier Valdivia et al. (2021); Wang et al. (2021a); Xu and Strohmer (2023). Prior work in causality has aimed to measure this tradeoff, including by quantifying the extent to which model loss increases when trained under causal fairness constraints Plečko and Bareinboim (2025) and by using the "average treatment effect" to assess the impact of fairness interventions Ji et al. (2024). However, the assumption that fairness and accuracy must exist as a "tradeoff" may be due to a lack of consideration of social and historical context (Wang et al., 2021a), limitations of the mathematical notations used to model fairness, and measurement error in biased data (Dutta et al., 2020). Our work explicitly considers social context and baseline bias in the observational data by accounting for non-direct causal effects, as observational health data is known to reflect the biases of our medical system. For a more in-depth treatment of the necessity of the fairness-accuracy tradeoff, we refer readers to Li and Li (2025). We emphasize the importance of using path-specific causal fairness to assess ML models in the healthcare setting, breaking down bias into direct, indirect, and spurious causal effects. This allows us to consider the model in the context of broader health disparities, quantify the extent to which a learned model replicates and/or exacerbates unfair causal paths in the data, and measure the impact of fairness interventions on model performance.
2.3 Fairness in foundation models
@@
Most work around fairness in foundation models focuses on language models (Wang et al., 2025; Gallegos et al., 2024) and vision models (Ali et al., 2023b). In medical imaging specifically, prior work found significant subgroup performance differences in segmentation foundation models Li et al. (2024); Jin et al. (2024). Other work echoes these findings and addresses these differences using balanced fine-tuning datasets Khan et al. (2023) or "fair PCA post-processing" Ali et al. (2023a). Consistent with broader ethical ML literature Chen et al. (2021), Queiroz et al. (2025) discusses the ways that bias mitigation for medical vision foundation models must be an integrated effort involving investment at all steps in the AI pipeline, from data collection to training methodologies to deployment and regulation. Relatively little work focuses on fairness in tabular foundation models. Robertson et al. (2025) demonstrates how training on causally fair synthetic data can improve the fairness of tabular foundation models. In the case of in-context learning, Kenfack et al. (2025) illustrates the promise of pre-processing fairness methods. Our work examines the path-specific causal fairness of structured (tabular) EHR foundation models, both at the pretraining step and after linear probing, which has received little attention in prior work.
3 Preliminaries
3.1 Standard Fairness Model
@@
The standard fairness model (SFM) (Plečko and Bareinboim, 2024) is a structural causal model (SCM) to assess causal fairness. An SCM (Pearl, 2000) is a formalized model of a data-generating process that defines the causal relationship between each of the variables in the system. The SFM defines the sensitive attribute ($X$) that can directly impact the outcome ($Y$). Additionally, $X$ affects $Y$ through a set of mediator variables ($W$), confounded by $Z$ (Figure 2a). One key aspect of the SFM is the identifiability of the natural direct effect (NDE), natural indirect effect (NIE), and spurious effect (SE) from observational data. That is, under the assumption that the data-generating mechanism follows the SFM, all three effects can be estimated as nested counterfactuals from observational data (Appendix 9).
3.2 Clinical mapping of Standard Fairness Model
@@
We leverage the SFM to define a generalizable procedure for mapping high-dimensional observational health data (e.g., EHRs, administrative claims) onto the SFM, thus providing a principled template for evaluating fairness in observational health settings. Such a generalizable mapping is important because relying on clinical domain knowledge to define structural assumptions is infeasible in high dimensions. Our mapping (Figure 2b) is as follows: as with the SFM, $X$ remains the sensitive attribute and $Y$ is the clinical outcome of interest. $W$ represents a high-dimensional representation of a person’s clinical record that can include information derived from diagnoses, medications, procedures, devices, laboratory tests, etc. These features can be derived directly from a person’s clinical record or can be a learned representation of the clinical data presented as a set of embedding vectors. Finally, $Z$ represents other demographic variables, baseline comorbidities known before a visit, if available, and healthcare utilization. Healthcare utilization (HCU) refers to the frequency, types, and patterns of healthcare services that an individual uses. HCU can serve as a proxy for several unobserved variables such as social determinants of health (typically constituting latent common causes between $X$ and $Z$), including access to and trust in the healthcare system. In our work, we operationalize healthcare utilization as the frequency of visits. The NDE represents the "direct" effect of the sensitive attribute on the outcome; the NIE represents the effect of the sensitive attribute on the outcome as mediated by the rest of a person’s observational health record ($W$). The SE represents the baseline change in the outcome distribution as a result of the two interventions $x=0$ and $x=1$ compared to their observational counterparts.
4 Methods
@@
Figure 3: Overview of the proposed workflow. Starting with task selection, we outline the steps required to train and evaluate a causally fair model. In particular, we find that data-driven decision making must be combined with known historical context about healthcare disparities specific to the clinical task at hand.
@@
Figure 4: Flowchart presenting the workflow outlined in this paper. The steps are: 1) define the task and key variables (in particular, select the sensitive attribute and calculate healthcare utilization); 2) robustly estimate the path-specific causal effects; 3) train a baseline model; 4) select target effects (e.g., NDE, NIE, SE, some combination) for which to reduce disparities; 5) perform and evaluate causal fairness enhancing interventions.
@@
4.1 Define the task and key variables
Once a clinical prediction task is selected, we use this to set our outcome ($Y$) in the SFM. The sensitive attribute ($X$), however, still requires definition. $X$ is typically a discrete variable, often binary, for which we want to measure and enforce path-specific fairness. While purely data-driven analysis based on all known sensitive attributes is important, it is often more crucial and feasible to assess how broadly known disparities reflect in the data and downstream learned models. We advocate for selecting a sensitive attribute that is associated with known (or hypothesized) disparities in broader health literature. We use race or gender as our main sensitive attributes for each case; however, selecting subgroups for $X$ using an intersection of identities is possible under statistical feasibility (sufficient sample sizes to robustly evaluate path-specific effects). The next key variable(s), often ignored, are related to healthcare utilization. Healthcare utilization patterns drive predictive performance of learned models  (Zink et al., 2024; Pang et al., 2025) and serve as noisy proxies for unobserved social drivers of health. We define healthcare utilization as the number of visits per year. If there is consistent information about the visit setting (e.g., outpatient, inpatient, pharmacy), we create separate healthcare utilization features for each visit setting. These healthcare utilization features, in combination with any "non-sensitive" demographic features, comprise the confounder set of the SFM ($Z$). $Z$ must also include baseline health status, such as known comorbidities known at the beginning of the visit or stay, if data samples represent visits or stays. The clinical features ($W$) include all features used by the model that are not encapsulated by $X$, $Z$, and $Y$.
4.2 Robustly estimate path-specific causal effects